Comlute the wavelength of the 9GHZ microwave in dielectric medium having dielectric constant 4
 

Grade:12th pass

1 Answers

Sahani Kumar
5 years ago
Given:
Frequency = 9 GHz = 9 * 10^9 Hz
We know that speed of light = 3 * 10^8 m/s
 
Wavelength =( speed of light)/(frequency) 
Wavelength=(3 *10^8)/(9 *10^9) = 1/30
Wavelength = 1/30 = 0.03
 
Regards 
Askiitian member 
If you like this answer please encourage us by approving this answer.